Using the following chemical equation how many liters of oxygen are consumed by burning 7.86 g of butane?

C4H10 + 5O2 -> 3CO2 + 4H2O


1
Expert's answer
2021-02-17T07:41:30-0500

C4H10 + 5O2 → 3CO2 + 4H2O - unbalanced equation


Solution:

C4H10 - butane

The balanced chemical equation:

2C4H10 + 13O2 → 8CO2 + 10H2O

According to the equation above: n(C4H10)/2 = n(O2)/13


Moles of C4H10 = Mass of C4H10 / Molar mass of C4H10

The molar mass of C4H10 is 58.12 g mol-1.

Hence,

n(C4H10) = 7.86 g / 58.12 g mol-1 = 0.1352 mol


n(O2) = 13 × n(C4H10) / 2 = 13 × 0.1352 mol / 2 = 0.8788 mol


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 0.8788 mol of oxygen (O2) occupies:

(0.8788 mol × 22.4 L) / 1 mol = 19.685 L of oxygen (O2)

0.8788 mol of oxygen (O2) occupies a volume equal to 19.69 L.


Answer: 19.69 liters of oxygen (O2) are consumed.


However, if an unbalanced equation is used, 15.15 liters of oxygen (O2) are consumed.